Calculate the H2O+ concentration for each pH. pH = 11 [H,0]= pH = 2 [H,0]= pH=7 [H,0]=
Answer #1

we know, pH = -log [#3 of 1 multiply with -2 a - pH = Log CH3 0] - apply log on both sides 10-PB = [H 3 87 Log (x) = lo* pe